# Majorana vortex phases in time-reversal invariant higher-order topological insulators and topologically trivial insulators

### v2

@@ I Introduction
Topological superconductors (TSCs) have emerged as a central topic in condensed matter physics for their ability to host Majorana zero modes (MZMs) [1, 2]. These quasiparticles, with their non-Abelian statistics, offer promising applications for topological quantum computing [3]. While definitive proof of TSCs remains elusive, substantial progress has been made in both theoretical [4, 5, 6, 7, 8, 9, 10, 11, 12, 13] and experimental [14, 15, 16, 17, 18, 19] realms over the past decade. A hallmark experimental signature of MZMs is the observation of zero-bias conductance peaks (ZBCPs) in tunneling spectroscopy measurements [20]. These ZBCPs have been reported within Abrikosov vortices across diverse systems, including proximity-induced superconducting topological insulator (TI) Bi2Te3/NbSe2 [15, 16], topological crystalline insulator (TCI) SnTe/Pb [19], and intrinsic superconductors FeTexSe1-x [21, 22, 23], (Li0.84Fe0.16)OHFeSe [24], 2M-WS2 [25, 26], and LiFeAs [27, 28, 29]. Central to understanding these observations is the presence of a gapless surface Dirac cone in the normal-state, as first proposed by Fu and Kane for superconducting TI [4]. Spin-momentum locking in the surface states induces an effective $p$-wave superconducting pairing, which enables the formation of MZMs at vortex cores [30]. Fu-Kane mechanism was also extended to Dirac and Weyl semimetals [31, 32, 33], which host gapless bulk Dirac and Weyl cones, respectively. Despite these advances, a key question remains: can Majorana vortex end modes (MVEMs) be realized in time-reversal invariant systems, which are compatible with superconducting pairing, in the absence of both bulk and surface gapless Dirac cones?
@@
Crystalline symmetries enrich the classification of topological phases, and their exploration leads to the discovery of higher-order topological insulators (HOTIs) [34, 35, 36, 37]. Over the past few years, higher-order topological phases, characterized by gapless modes at boundaries with a codimension greater than one, have attracted significant research interest due to their distinctive bulk-boundary correspondence [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59]. A prominent example is time-reversal-invariant HOTIs (THOTIs) in three dimensions [60, 61, 62], which host gapped surfaces but feature robust helical hinge modes (HHMs). THOTIs can be realized in double band inversions systems with additional crystalline symmetries, such as rotation [60, 37] or inversion [63, 64, 65, 66]. By using symmetry indicator theory and first-principles calculations, THOTIs have been theoretically predicted in a diverse range of materials [67, 68, 69, 70]. Recently, significant experimental progress has been made in realizing THOTIs, with evidence of HHMs reported in bismuth (Bi) [71, 72, 73, 74, 75, 76], BixSb1-x [74], 1T-XTe2 (X=Mo,W) [77, 78, 79], and Bi4X4 (X=Br, I)[80, 81, 82, 83, 84, 85, 86]. In particular, Bi4X4 was predicted to exhibit superconductivity under pressure [87, 88], and experiments have demonstrated superconducting proximity effects in Bi [89, 90, 91, 92, 93, 94].
@@
Motivated by recent experimental advances in superconducting THOTIs, we investigate Majorana vortex phases in these systems. We construct THOTIs by combining two decoupled TIs (double TIs) [Fig. 1(a)] and time-reversal symmetry-preserving mass terms that anisotropically gap the surface states, enabling the emergence of HHMs. This approach enables the realization of both intrinsic THOTIs [38], protected by rotational [Fig. 1(b)] or inversion [Fig. 1(d)] symmetry, and extrinsic THOTI [Fig. 1(c)] with trivial bulk topology [38]. In the absence of additional mass terms, the double TIs system undergoes two distinct vortex phase transitions at critical chemical potentials $\mu_{c}^{(1)}$ and $\mu_{c}^{(2)}$. Remarkably, the introduction of mass terms has a negligible impact on these transitions and does not induce additional topological phase transitions [Figs. 1(e)-1(h)]. Consequently, MVEMs emerge in THOTI systems when the chemical potential lies within the range $\mu_{c}^{(1)}<\mu<\mu_{c}^{(2)}$ [Fig. 2(b)] due to the the $Z_{2}$ classification for systems within the D symmetry class. We further confirm this behavior in rotational and inversion symmetries-protected THOTI Bi [71] through effective lattice model analysis, observing similar conditions for MVEMs emergence. Remarkably, MVEMs persist even when all surfaces of double TIs or TI are gapped with the same signs, rendering the system topologically trivial in both first- and second-order categories [Fig. 4]. These models provide paradigms for realizing Majorana vortex phases in topologically trivial systems.
@@ II Construction of THOTIs
TIs with inversion symmetry are characterized by $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ invariants, determined by inversion eigenvalues at time-reversal invariant momenta [95]. In contrast, symmetry indicator theory predicts that time-reversal invariant systems possessing inversion symmetry are described by a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ invariant structure [67]. This discrepancy reveals a THOTI phase characterized by a $\mathbb{Z}_{4}$ symmetry indicator invariant $\kappa=2$, arising from double band inversions at time-reversal invariant points. Additionally, rotational symmetry-protected TCIs are predicted to host gapless Dirac cones on surfaces perpendicular to an $n$-fold rotational axis, which are connected by $n$ one-dimensional (1D) HHMs along hinges parallel to the rotation axis [37]. Consequently, such TCIs can also be classified as THOTIs due to the presence of HHMs. Using a combination of symmetry indicator analysis and first-principles calculations, THOTIs protected by either inversion or rotational symmetry have been predicted in a wide range of materials [67, 68, 69, 70]. In the following, we construct intrinsic THOTIs stabilized by these crystalline symmetries, as well as extrinsic THOTIs that lack crystalline symmetry protection.
@@
When $\eta_{0}=0$, the higher-order topology of $H_{\mathrm{HOTI}}^{(1)}$ was studied in Ref. 60. This THOTI is protected by the time-reversal symmetry and four-fold rotational symmetry around the $z$-axis, represented by $C_{4z}=\tau_{z}e^{-i\pi s_{z}/4}$. In contrast, $H_{\mathrm{HOTI}}^{(2)}$ breaks the $C_{4z}$ symmetry but preserves the inversion symmetry, represented by $I=\tau_{0}\sigma_{z}s_{0}$. With $T$ and $I$ symmetries, $H_{\mathrm{HOTI}}^{(2)}$ has a $\mathbb{Z}_{4}$ classification and $\kappa=2$ for this symmetry indicator invariant because of the double band inversions at the $\Gamma$ point. This indicates that $H_{\mathrm{HOTI}}^{(2)}$ describes a inversion symmetry-protected THOTI [64].
The non-trivial higher-order topology of $H_{\mathrm{HOTI}}^{(1)}$ can be revealed from its surface state analysis. For $\eta(\bm{k})=0$, $H_{\mathrm{HOTI}}^{(1)}$ decouples into two copies of TIs. In this configuration, the system exhibits two gapless Dirac cones on each of the (100), (010), and (001) surfaces. Introducing a non-zero $\eta_{1}$ while maintaining $\eta_{0}=0$, the two Dirac cones on the (100) and (010) surfaces are gapped with opposite mass signs, enforced by the $C_{4z}$ symmetry. This results in gapless HHMs along the $z$-direction. Since $\eta(\bm{k})=0$ at the $\Gamma$ point, the (001) surface remains gapless, hosting two Dirac cones. A schematic illustration of this $C_{4z}$ symmetry-protected THOTI is presented in Fig. 1(b). When $\eta_{0}$ is non-zero, the $C_{4z}$ symmetry is broken, leading to a gapped (001) surface with a gap magnitude of $\eta_{0}$. In this scenario, the two Dirac cones on the (100) and (010) surfaces acquire gaps of magnitudes $\eta_{0}-\eta_{1}/2$ and $\eta_{0}+\eta_{1}/2$, respectively (see Appendix A). Consequently, for $20<\eta_{0}<\eta_{1}/$ 2, the (100) surface develops a gap with a sign opposite to that of the (010) and (001) surfaces. This configuration results in HHMs along the edges of the (100) surface, as schematically illustrated in Fig. 1(c). Although the absence of crystalline symmetry protection, this extrinsic THOTI phase is protected by the time-reversal symmetry and surface energy gaps, ensuring the robustness of HHMs [38].
@@ III Majorana vortex phases in THOTIs
Notably, although the superconducting THOTI systems in cases (ii)-(iv) exhibit identical vortex phase transitions, the TSC phases for case (ii) differ from those in cases (iii) and (iv). In case (ii), the system hosts the $\mathcal{C}_{4z}$ symmetry, which enriches the topological classification from $\mathbb{Z}_{2}$ to $\mathbb{Z}_{2}\otimes\mathbb{Z}_{2}$ [97]. In cases (iii) and (iv), the system hosts the $\mathbb{Z}_{2}$ classification within the D symmetry class. Specifically, in case (ii), the two vortex phase transitions are distinguished by the $\mathcal{C}_{4z}$ symmetry eigenvalues, as indicated by the color coding in Fig. 1(f). This implies that, for $\mu<\mu_{c}^{(1)}$, both the Hamiltonians, in the subsectors having the $\mathcal{C}_{4z}$ symmetry eigenvalues $1$ and $-1$, are topologically nontrivial, and the two subsectors cannot be coupled because of the $\mathcal{C}_{4z}$ symmetry. Consequently, the number of robust MVEMs changes as $2\rightarrow 1\rightarrow 0$ when $\mu$ crosses ${\mu}_{c}^{(1)}$ and ${\mu}_{c}^{(2)}$ in sequence, as schematically depicted in Fig. 2(a). In contrast, for cases (iii) and (iv), no crystalline symmetries can protect two MVEMs. Consequently, MVEMs are present on the (001) surface only within the region ${\mu}_{c}^{(1)}<\mu<{\mu}_{c}^{(2)}$, as schematically illustrated in Fig. 2(b). This scenario is also applicable to the side surfaces of systems in cases (ii)-(iv), owing to the absence of rotational symmetry for these surfaces. It is noted that when $\lambda=0$, ${\mu}_{c}^{(1)}={\mu}_{c}^{(2)}$, and therefore no $\mathbb{Z}_{2}$-protected MVEMs exist.
@@ IV MVEMs in superconducting bismuth
Bismuth-based materials have played a pivotal role in the advancement of topological physics [99, 100]. However, pure bismuth, exhibiting two band inversions [101], was long been thought to be topologically trivial due to the $Z_{2}$ classification for strong TI. The discovery of higher-order topological phases has reclassified bismuth as a THOTI, protected by inversion symmetry and three-fold rotational symmetry $C_{3z}$ along its trigonal axis [71]. The higher-order topology of bismuth has been robustly validated through multiple experimental observations by the scanning tunneling microscopy measurements [71, 73, 74, 75, 94] and Josephson junction studies [71, 72, 76]. Recently, the superconducting proximity effect in bismuth was demonstrated [94], and ZBCP, indicative of MZMs, was observed at the interface between a superconductor and an Fe cluster at the edge of bismuth [90]. Especially, ZBCPs were observed for Bi islands on superconductors Fe(Te,Se) [91, 92]. In the following, we employ the effective tight-binding model developed in Ref. 71, which captures the higher-order topology of bismuth, to investigate the Majorana vortex phase.
@@
where $H_{\mathrm{TB,I}}(\bm{k})$ and $H_{\mathrm{TB,II}}(\bm{k})$ describe a strong TI within distinct eigenvalue sectors of the $C_{3z}$ symmetry. Specifically, $H_{\mathrm{TB,I}}$ corresponds to $C_{3z}$ eigenvalues $\exp(\pm i\pi/3)$, and $H_{\mathrm{TB,II}}$ corresponds to an eigenvalue of $-1$. The $C_{3z}$ symmetry operator is expressed as $C_{3z}=C_{3,\mathrm{I}}^{z}\oplus C_{3,\mathrm{II}}^{z}$, with $C_{3,\mathrm{I}}^{z}=\sigma_{0}\otimes e^{i\frac{\pi}{3}\sigma_{3}}$ and $C_{3,\mathrm{II}}^{z}=-\sigma_{0}\otimes\sigma_{0}$. The term $M_{\mathrm{TB}}(\bm{k})$ describes the coupling between $H_{\mathrm{TB,I}}$ and $H_{\mathrm{TB,II}}$, with the coupling strength controlled by the parameter $\delta$. The parameter $\epsilon$ introduces a relative energy shift between $H_{\mathrm{TB,I}}$ and $H_{\mathrm{TB,II}}$. Detailed expressions for $H_{\mathrm{TB,I}}(\bm{k})$, $H_{\mathrm{TB,II}}(\bm{k})$, and $M_{\mathrm{TB}}(\bm{k})$, along with specific model parameters, are provided in Appendix D. When $\epsilon=0$, the surface states of bismuth feature two Dirac points, each contributed by $H_{\mathrm{TB,I}}$ and $H_{\mathrm{TB,II}}$, respectively, and pinned at the same energy. In this scenario, the two Dirac cones cannot be gapped while preserving both $C_{3z}$ symmetry (preventing inter-cone gapping) and time-reversal symmetry (preventing intra-cone gapping). However, no symmetry constraints fix the energies of the two Dirac cones, allowing them to be gapped for $\epsilon\neq 0$ [71]. Enforced by the inversion and $C_{3z}$ symmetries, the surface states related by inversion exhibit opposite mass signs, while those related by three-fold rotation share identical mass signs. This energy gap configuration of the surface states gives rise to HHMs along specific direction, as schematically illustrated in Fig. 3(a).
By replacing $H_{\text{HOTI}}$ with $H_{\text{TB}}$ in Eq. (2), we diagonalize the BdG Hamiltonian $\mathcal{H}_{\text{BdG}}(x,y,k_{z}=\pi)$ and identify two vortex phase transitions at chemical potentials $\mu_{c}^{(3)}$ and $\mu_{c}^{(4)}$, respectively. These transitions arise at $k_{z}=\pi$ due to double band inversions at the $T$ point [$\bm{k}=(0,0,\pi)$] under the selected model parameters [71]. This result aligns with the analysis in Sec. III. Unlike the $C_{4z}$ symmetry, the $C_{3z}$ symmetry in bismuth does not change the classification of MVEMs [97, 98], hoting the $\mathbb{Z}_{2}$ topological classification. Consequently, a single MZM emerges at the vortex core when the chemical potential lies within $\mu_{c}^{(3)}<\mu<\mu_{c}^{(4)}$.

### v3

@@ I Introduction
Topological superconductors (TSCs) have emerged as a central topic in condensed matter physics due to their ability to host Majorana zero modes (MZMs) [20, 64]. These quasiparticles, with their non-Abelian statistics, offer promising applications for topological quantum computing [55]. Although definitive proof of MZMs remains elusive, substantial progress has been made in both theoretical [17, 67, 63, 60, 53, 85, 61, 57, 48, 98] and experimental [51, 86, 73, 94, 15, 40] realms over the past decade. A hallmark experimental signature of MZMs is the observation of zero-bias conductance peaks (ZBCPs) in tunneling spectroscopy measurements [33]. These ZBCPs have been reported within Abrikosov vortices across diverse systems, including proximity-induced superconducting topological insulator (TI) Bi2Te3/NbSe2 [86, 73], topological crystalline insulator (TCI) SnTe/Pb [40], and intrinsic superconductors FeTexSe1-x [78, 106, 30], (Li0.84Fe0.16)OHFeSe [39], 2M-WS2 [92, 13], and LiFeAs [29, 41, 36]. Central to understanding these observations is the presence of a gapless surface Dirac cone in the normal-state, as first proposed by Fu and Kane for superconducting TI [17]. Spin-momentum locking in the surface states induces an effective $p$-wave superconducting pairing, which enables the formation of MZMs at vortex cores [24]. Fu-Kane mechanism was also extended to Dirac and Weyl semimetals [66, 89, 97], which host gapless bulk Dirac and Weyl cones, respectively. Despite these advances, a key question remains: can Majorana vortex end modes (MVEMs) be realized in time-reversal invariant systems, which are compatible with superconducting pairing, in the absence of both bulk and surface gapless Dirac cones?
@@
Crystalline symmetries enrich the classification of topological phases, and their exploration leads to the discovery of higher-order topological insulators (HOTIs) [3, 2, 32, 14]. Over the past few years, higher-order topological phases, characterized by gapless modes at boundaries with a codimension greater than one, have attracted significant research interest due to their distinctive bulk-boundary correspondence [18, 88, 79, 59, 95, 96, 87, 93, 70, 82, 7, 46, 58, 99, 45, 49, 35, 43, 50, 101, 100, 44]. A prominent class is time-reversal-invariant HOTIs (THOTIs) in three dimensions [72, 68, 90], which host gapped surfaces but feature robust helical hinge modes (HHMs). THOTIs can be realized in double band inversions systems with additional crystalline symmetries, such as rotation [72, 14] or inversion [62, 27, 26, 80]. By using symmetry indicator theory and first-principles calculations, THOTIs have been theoretically predicted in a diverse range of materials [75, 74, 77, 5]. Recently, significant experimental progress has been made in realizing THOTIs, with evidence of HHMs reported in bismuth (Bi) [69, 52, 25, 54, 1, 4], BixSb1-x [54], 1T-XTe2 (X=Mo,W) [31, 11, 34], and Bi4X4 (X=Br, I)[23, 56, 71, 104, 103, 91, 105]. In particular, Bi4X4 was predicted to exhibit superconductivity under pressure [65, 37], and experiments have demonstrated superconducting proximity effects in Bi [12, BertholdJäck2019, 8, 6, 38, 102].
@@
Motivated by recent experimental advances in superconducting THOTIs, we investigate Majorana vortex phases in these systems. We construct THOTIs by combining two decoupled TIs (double TIs) [Fig. 1(a)] and time-reversal symmetry-preserving mass terms that anisotropically gap the surface states, enabling the emergence of HHMs. This approach enables the realization of both intrinsic THOTIs [18], protected by rotational [Fig. 1(b)] or inversion [Fig. 1(d)] symmetry, and extrinsic THOTI [Fig. 1(c)] with trivial bulk topology [18]. In the absence of additional mass terms, the double TIs system undergoes two distinct vortex phase transitions at critical chemical potentials $\mu_{c}^{(1)}$ and $\mu_{c}^{(2)}$. Remarkably, the introduction of mass terms has a negligible impact on these transitions and does not induce additional topological phase transitions [Figs. 1(e)-1(h)]. Consequently, MVEMs emerge in THOTI systems when the chemical potential lies within the range $\mu_{c}^{(1)}<\mu<\mu_{c}^{(2)}$ [Fig. 2(b)] due to the $Z_{2}$ classification for systems within the D symmetry class. We further confirm this behavior in rotational and inversion symmetries-protected THOTI Bi [69] through effective lattice model analysis, observing similar conditions for the emergence of MVEMs . Remarkably, MVEMs persist even when all surfaces of double TIs or TI are gapped with the same signs, rendering the system topologically trivial in both first- and second-order categories [Fig. 4]. These models provide paradigms for realizing Majorana vortex phases in topologically trivial systems.
@@ II Construction of THOTIs
TIs are characterized by $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ invariants, which can be determined by inversion eigenvalues at time-reversal invariant momenta if inversion symmetry is preserved [16]. However, symmetry indicator theory predicts that time-reversal invariant systems possessing inversion symmetry are described by a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ invariant structure [75]. This discrepancy reveals a THOTI phase characterized by a $\mathbb{Z}_{4}$ symmetry indicator invariant $\kappa=2$, arising from double band inversions at time-reversal invariant points. Additionally, rotational symmetry-protected TCIs are predicted to host gapless Dirac cones on surfaces perpendicular to an $n$-fold rotational axis, which are connected by $n$ one-dimensional (1D) HHMs along hinges parallel to the rotation axis [14]. Consequently, such TCIs are also classified as THOTIs due to the presence of HHMs. Using a combination of symmetry indicator analysis and first-principles calculations, THOTIs protected by either inversion or rotational symmetry have been predicted in a wide range of materials [75, 74, 77, 5]. In the following, we construct intrinsic THOTIs stabilized by these crystalline symmetries, as well as extrinsic THOTIs that lack crystalline symmetry protection.
@@
When $\eta_{0}=0$, the higher-order topology of $H_{\mathrm{HOTI}}^{(1)}$ was studied in Ref. 72. This THOTI is protected by the time-reversal symmetry and four-fold rotational symmetry around the $z$-axis, represented by $C_{4z}=\tau_{z}e^{-i\pi s_{z}/4}$. In contrast, $H_{\mathrm{HOTI}}^{(2)}$ breaks the $C_{4z}$ symmetry but preserves the inversion symmetry, represented by $I=\tau_{0}\sigma_{z}s_{0}$. With $T$ and $I$ symmetries, $H_{\mathrm{HOTI}}^{(2)}$ has a $\mathbb{Z}_{4}$ classification and $\kappa=2$ for this symmetry indicator invariant because of the double band inversions at the $\Gamma$ point. This indicates that $H_{\mathrm{HOTI}}^{(2)}$ describes a inversion symmetry-protected THOTI [27].
The non-trivial higher-order topology of $H_{\mathrm{HOTI}}^{(1)}$ can be revealed from its surface state analysis. For $\eta(\bm{k})=0$, $H_{\mathrm{HOTI}}^{(1)}$ decouples into two copies of TIs. In this configuration, the system exhibits two gapless Dirac cones on each of the (100), (010), and (001) surfaces. Introducing a non-zero $\eta_{1}$ while maintaining $\eta_{0}=0$, the two Dirac cones on the (100) and (010) surfaces are gapped with opposite mass signs, enforced by the $C_{4z}$ symmetry. This results in gapless HHMs along the $z$-direction. Since $\eta(\bm{k})=0$ at the $\Gamma$ point, the (001) surface remains gapless, hosting two Dirac cones. A schematic illustration of this $C_{4z}$ symmetry-protected THOTI is presented in Fig. 1(b). When $\eta_{0}$ is non-zero, the $C_{4z}$ symmetry is broken, leading to a gapped (001) surface with a gap magnitude of $\eta_{0}$. In this scenario, the two Dirac cones on the (100) and (010) surfaces acquire gaps of magnitudes $\eta_{0}-\eta_{1}/2$ and $\eta_{0}+\eta_{1}/2$, respectively (see Appendix A). Consequently, for $0<\eta_{0}<\eta_{1}/2$, the (100) surface develops a gap with a sign opposite to that of the (010) and (001) surfaces. This configuration results in HHMs along the edges of the (100) surface, as schematically illustrated in Fig. 1(c). Although the absence of crystalline symmetry protection, this extrinsic THOTI phase is protected by the time-reversal symmetry and surface energy gaps, ensuring the robustness of HHMs [18].
@@ III Majorana vortex phases in THOTIs
Notably, although the superconducting THOTI systems in cases (ii)-(iv) exhibit identical vortex phase transitions, the TSC phases for case (ii) differ from those in cases (iii) and (iv). In case (ii), the system hosts the $\mathcal{C}_{4z}$ symmetry, which enriches the topological classification from $\mathbb{Z}_{2}$ to $\mathbb{Z}_{2}\otimes\mathbb{Z}_{2}$ [28]. In cases (iii) and (iv), the system hosts the $\mathbb{Z}_{2}$ classification within the D symmetry class. Specifically, in case (ii), the two vortex phase transitions are distinguished by the $\mathcal{C}_{4z}$ symmetry eigenvalues, as indicated by the color coding in Fig. 1(f). This implies that, for $\mu<\mu_{c}^{(1)}$, both the Hamiltonians, in the subsectors having the $\mathcal{C}_{4z}$ symmetry eigenvalues $1$ and 1-1, are topologically nontrivial, and the two subsectors cannot be coupled because of the $\mathcal{C}_{4z}$ symmetry. Consequently, the number of robust MVEMs changes as $2\rightarrow 1\rightarrow 0$ when $\mu$ crosses ${\mu}_{c}^{(1)}$ and ${\mu}_{c}^{(2)}$ in sequence, as schematically depicted in Fig. 2(a). In contrast, for cases (iii) and (iv), no crystalline symmetries can protect two MVEMs. Consequently, MVEMs are present on the (001) surface only within the region ${\mu}_{c}^{(1)}<\mu<{\mu}_{c}^{(2)}$, as schematically illustrated in Fig. 2(b). This scenario is also applicable to the side surfaces of systems in cases (ii)-(iv), owing to the absence of rotational symmetry for these surfaces. It is noted that when $\lambda=0$, ${\mu}_{c}^{(1)}={\mu}_{c}^{(2)}$, and therefore no $\mathbb{Z}_{2}$-protected MVEMs exist.
@@ IV MVEMs in superconducting bismuth
Bismuth-based materials have played a pivotal role in the advancement of topological physics [84, 42]. However, pure bismuth, exhibiting two band inversions [76], was long been thought to be topologically trivial due to the $Z_{2}$ classification for strong TI. The discovery of higher-order topological phases has reclassified bismuth as a THOTI, protected by inversion symmetry and three-fold rotational symmetry $C_{3z}$ along its trigonal axis [69]. The higher-order topology of bismuth has been robustly validated through multiple experimental observations by the scanning tunneling microscopy measurements [69, 25, 54, 1, 102] and Josephson junction studies [69, 52, 4]. Recently, the superconducting proximity effect in bismuth was demonstrated [102], and ZBCP, indicative of MZMs, was observed at the interface between a superconductor and an Fe cluster at the edge of bismuth [BertholdJäck2019]. Especially, ZBCPs were observed for Bi islands on superconductors Fe(Te,Se) [8, 6]. In the following, we employ the effective tight-binding model developed in Ref. 69, which captures the higher-order topology of bismuth, to investigate the Majorana vortex phase.
@@
where $H_{\mathrm{TB,I}}(\bm{k})$ and $H_{\mathrm{TB,II}}(\bm{k})$ describe a strong TI within distinct eigenvalue sectors of the $C_{3z}$ symmetry. Specifically, $H_{\mathrm{TB,I}}$ corresponds to $C_{3z}$ eigenvalues $\exp(\pm i\pi/3)$, and $H_{\mathrm{TB,II}}$ corresponds to an eigenvalue of 1-1. The $C_{3z}$ symmetry operator is expressed as $C_{3z}=C_{3,\mathrm{I}}^{z}\oplus C_{3,\mathrm{II}}^{z}$, with $C_{3,\mathrm{I}}^{z}=\sigma_{0}\otimes e^{i\frac{\pi}{3}\sigma_{3}}$ and $C_{3,\mathrm{II}}^{z}=-\sigma_{0}\otimes\sigma_{0}$. The term $M_{\mathrm{TB}}(\bm{k})$ describes the coupling between $H_{\mathrm{TB,I}}$ and $H_{\mathrm{TB,II}}$, with the coupling strength controlled by the parameter $\delta$. The parameter $\epsilon$ introduces a relative energy shift between $H_{\mathrm{TB,I}}$ and $H_{\mathrm{TB,II}}$. Detailed expressions for $H_{\mathrm{TB,I}}(\bm{k})$, $H_{\mathrm{TB,II}}(\bm{k})$, and $M_{\mathrm{TB}}(\bm{k})$, along with specific model parameters, are provided in Appendix D. When $\epsilon=0$, the surface states of bismuth feature two Dirac points, each contributed by $H_{\mathrm{TB,I}}$ and $H_{\mathrm{TB,II}}$, respectively, and pinned at the same energy. In this scenario, the two Dirac cones cannot be gapped while preserving both $C_{3z}$ symmetry (preventing inter-cone gapping) and time-reversal symmetry (preventing intra-cone gapping). However, no symmetry constraints fix the energies of the two Dirac cones, allowing them to be gapped for $\epsilon\neq 0$ [69]. Enforced by the inversion and $C_{3z}$ symmetries, the surface states related by inversion exhibit opposite mass signs, while those related by three-fold rotation share identical mass signs. This energy gap configuration of the surface states gives rise to HHMs along specific direction, as schematically illustrated in Fig. 3(a).
By replacing $H_{\text{HOTI}}$ with $H_{\text{TB}}$ in Eq. (2), we diagonalize the BdG Hamiltonian $\mathcal{H}_{\text{BdG}}(x,y,k_{z}=\pi)$ and identify two vortex phase transitions at chemical potentials $\mu_{c}^{(3)}$ and $\mu_{c}^{(4)}$, respectively. These transitions arise at $k_{z}=\pi$ due to double band inversions at the $T$ point $\bm{k}=(0,0,\pi)$] under the selected model parameters [69]. This result aligns with the analysis in Sec. III. Unlike the $C_{4z}$ symmetry, the $C_{3z}$ symmetry in bismuth does not change the classification of MVEMs [28, 47], hosting the $\mathbb{Z}_{2}$ topological classification. Consequently, a single MZM emerges at the vortex core when the chemical potential lies within $\mu_{c}^{(3)}<\mu<\mu_{c}^{(4)}$.